OpenStudy (anonymous):

Which of the following is equivalent to: (4z+12)^4

OpenStudy (anonymous):

256(z + 3)^4 4(z + 12)^4 4(z + 3)^4 256(z + 12)^4

OpenStudy (anonymous):

I need to know how to do this problem

mathslover (mathslover):

\[\large{(4z+12)^4 = [(4)(z+3)]^4}\]

mathslover (mathslover):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but thats not the right answer

OpenStudy (anonymous):

256(z + 3)^4 <==== is the answer

mathslover (mathslover):

wait : \[\large{4^4 (z+3)^4 = 256 (z+3)^4}\]

mathslover (mathslover):

4^4 = 256

mathslover (mathslover):

hence the answer is : \[\large{256(z+3)^4}\]

mathslover (mathslover):

got it @Brent0423 ?

OpenStudy (anonymous):

could you further simplify it

mathslover (mathslover):

Yes but there is no need of it

mathslover (mathslover):

Since the options need no simplification further .. r u getting my point @Brent0423 ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what about (x-5)^3

mathslover (mathslover):

\[\large{(x-5)^3=(x-5)(x-5)(x-5)}\]

mathslover (mathslover):

\[\large{(x-5)^2(x-5)=(x-5)^3}\] right?

OpenStudy (anonymous):

ya

mathslover (mathslover):

what is (x-5)^2 ?

OpenStudy (anonymous):

(x-5)(x-5) x^2-10x+25

mathslover (mathslover):

right good work , now we have: \[\large{(x-5)^2(x-5)=(x^2-10x+25)(x-5)}\]

mathslover (mathslover):

\[\large{(x^2-10x+25)(x-5)}\] \[\large{(x^2)(x)-5x^2-10x(x)-10x(-5)+25x-25(5)}\]

OpenStudy (anonymous):

okay i got it!

mathslover (mathslover):

good you can multiply them and solve.. also : the formula for (a-b)^3 is: \[\large{(a-b)^3=(a+b)(a^2-ab+b^2)}\]

OpenStudy (anonymous):

X^3-15x^2+75x-125

mathslover (mathslover):

good gr8 work

OpenStudy (anonymous):

Which of the following is equivalent to \[\frac{ \frac{ t+A }{ 5 } }{ t+A }, t \neq-A\]

OpenStudy (anonymous):

how do i figure this one out?

mathslover (mathslover):

ask this as a new question .. it will be answered there..